### v1

# BrainMorph: A Foundational Keypoint Model for Robust and Flexible Brain MRI Registration

Abstract
We present a keypoint-based foundation model for general purpose brain MRI registration, based on the recently-proposed KeyMorph framework. Our model, called BrainMorph, serves as a tool that supports multi-modal, pairwise, and scalable groupwise registration. BrainMorph is trained on a massive dataset of over 100,000 3D volumes, skull-stripped and non-skull-stripped, from nearly 16,000 unique healthy and diseased subjects. BrainMorph is robust to large misalignments, interpretable via interrogating automatically-extracted keypoints, and enables rapid and controllable generation of many plausible transformations with different alignment types and different degrees of nonlinearity at test-time. We demonstrate the superiority of BrainMorph in solving 3D rigid, affine, and nonlinear registration on a variety of multi-modal brain MRI scans of healthy and diseased subjects, in both the pairwise and groupwise setting. In particular, we show registration accuracy and speeds that surpass current state-of-the-art methods, especially in the context of large initial misalignments and large group settings. All code and models are available at https://github.com/alanqrwang/brainmorph.
{IEEEkeywords}
Image registration, Multi-modal, Keypoint detection, Foundation model, Brain MRI
1 Introduction
[…]
Registration is a fundamental problem in biomedical imaging tasks. Multiple images, often reflecting a variety of contrasts, modalities, subjects, and underlying pathologies, are commonly acquired in many applications [1]. Registration seeks to spatially align these images in order to facilitate downstream analyses, like tracking longitudinal changes, studying disease progression, or analyzing population-level variability.
Registration can be broken down into different types. It may be performed within the same modality (unimodal) or across different modalities (multimodal). Pairwise registration performs registration on an image pair, while groupwise registration performs registration on multiple images at once [2]. To perform the registration, various families of spatial transformations can be used, including rigid, affine, and nonlinear transformations.
Different lines of research have been explored to solve the registration task. “Classical” (i.e. non-learning-based) registration methods solve an iterative optimization of a similarity metric over a space of transformations, with additional regularization terms that restrict the space of plausible transformations [3, 4]. Much research is dedicated to developing good transformations, similarity metrics, and optimization strategies. While performant, these approaches are known to suffer from long run times, often requiring upwards of several minutes to register a pair of images. In addition, these approaches are known to perform poorly when the initial misalignment between images is large (e.g. 90 degrees of rotational misalignment).
Table 1: Summary of registration capabilities of BrainMorph compared to baselines.
Support for… ITK-Elastix ANTs SynthMorph BrainMorph
pairwise?
groupwise?
multi-modal?
rigid?
affine?
nonlinear?
lesions?
non-skullstripped?
Another line of work decomposes the registration problem into two steps. First, salient features (e.g. keypoints or contours) are extracted from images, and correspondences are established between the features of the image pair. Second, the transformation is derived which aligns these features and correspondences. In this work, we refer to methods that extract keypoints as salient features as “keypoint-based” registration. Largely, keypoint-based registration is advantageous in that the registration is relatively robust to initial misalignments, given good correspondences. In addition, these methods enjoy superior interpretability, because the user can interrogate the correspondences which are driving the registration. However, finding keypoints and establishing correspondences is a difficult task and is a subject of much research [5, 6, 7].
More recently, deep learning-based strategies have emerged which leverage large datasets of images to train a neural network to perform the registration task. These strategies use convolutional neural network (CNN) or transformer-based architectures [8, 9, 10] that either output transformation parameters (e.g. rigid, affine or spline) [11, 12] or a dense deformation field [13, 14] which aligns an image pair. These strategies are effective and are able to perform fast inference via efficient feed-forward passes. However, like classical methods, they often fail when the initial misalignment is large.
Figure 1: The framework of BrainMorph. Fixed and moving 3D brain images are passed through the same keypoint detection network, which predicts $N$ keypoints useful for registration. The transformation parameters are then computed as a function of the keypoints, which are in turn used to resample the moving image. Keypoint colors denote depth (see Fig. 2).
Recently, the KeyMorph framework proposed to combine the benefits of keypoint-based registration with deep learning using neural networks to automatically detect corresponding keypoints [15, 16]. Corresponding keypoints can then be used to compute the optimal transformation in closed-form, where the keypoints themselves are learned by a neural network. Thus, KeyMorph may be seen as possessing all the benefits of keypoint-based registration, including robustness to large misalignments and better interpretability, while retaining the fast inference times of deep learning-based methods. In addition, different transformation can be used according to user specifications, thereby enabling human controllability of the registration process.
In this work, we extend the KeyMorph framework into a general-purpose tool for brain MRI registration. Although the literature on deep-learning-based registration methods has been widely explored and many tools exist with support for brain MRIs [17, 18, 14], most works focus on cross-subject pairwise registration on healthy subjects with skull-stripped images. Groupwise registration and support for non-skull-stripped and diseased subjects is often ignored, or is prohibitively slow or memory-intensive [18]. Often, the tools that do exist require time-consuming preprocessing steps like skullstripping and pre-affine registration [13, 14]. In short, there is a lack of software tools for brain MRIs which is capable of supporting registration across a wide swath of use cases, including healthy and diseased subjects, pairwise/longitudinal/groupwise registration, and minimal assumptions on preprocessing (like skullstripping).
Table 1 summarizes the capabilities of our tool, which we call BrainMorph. Our tool supports rigid, affine, and nonlinear registration. It is trained, and thus works “out-of-the-box”, on both skull-stripped and non-skull-stripped data of diseased and normal subjects in a variety of MRI modalities. In addition to pairwise registration, we introduce a novel and memory-efficient approach to groupwise registration and demonstrate the superiority and scalability of our approach in population-level and longitudinal settings. All code and models are available at https://github.com/alanqrwang/brainmorph.
2 Background
Classical Methods. Pairwise iterative, optimization-based approaches have been extensively studied in medical image registration [19, 3]. These methods employ a variety of similarity functions, types of deformation, transformation constraints or regularization strategies, and optimization techniques. Intensity-based similarity criteria are most often used, such as mean-squared error (MSE) or normalized cross correlation for registering images of the same modality [20, 21, 22]. For registering image pairs from different modalities, statistical measures like mutual information or contrast-invariant features like MIND are popular [23, 22, 24, 25, 26].
Keypoint-based Methods. Another registration paradigm first detects features or keypoints in the images, and then establishes their correspondence. This approach often involves handcrafted features [27], features extracted from curvature of contours [28], image intensity [29, 30], color information [31, 32], or segmented regions [33, 34]. Features can be also obtained so that they are invariant to viewpoints [35, 36, 5, 37]. These algorithms then optimize similarity functions based on these features over the space of transformations [38, 19]. This strategy is sensitive to the quality of the keypoints and often suffer in the presence of substantial contrast and/or color variation [39].
Deep Learning-based Methods. In learning-based image registration, supervision can be provided through ground-truth transformations, either synthesized or computed by classical methods [40, 41, 42, 43, 44, 45]. Unsupervised strategies use loss functions similar to those employed in classical methods [13, 46, 12, 47, 48, 49, 50, 51]. Weakly supervised models employ (additional) landmarks or labels to guide training [13, 52, 53, 54].
Recent learning-based methods compute image features or keypoints [55, 56] that can be used for image recognition, retrieval, tracking, or registration. Learning useful features or keypoints can be done with supervision [39, 57, 58], self-supervision [59, 60] or without supervision [61, 62, 63]. Finding correspondences between pairs of images usually involves identifying the learned features which are most similar between the pair. In contrast, our method uses a network which extract/generates keypoints directly from the image. The keypoints between the moving and fixed image are corresponding (i.e., matched) by construction, and we optimize these corresponding keypoints directly for the registration task (and not using any intermediate keypoint supervision).
Learning-based methods for multi-modal registration are of great practical utility and often-studied in the literature. Most works require, in addition to the moving and fixed image, a corresponding image in a standard space which can be compared and which drives the alignment, usually in the form of segmentations. [64] address multi-modal retinal images and handle multi-modality by transforming each image to a standard grayscale image via vessel segmentation. A standard feature detection and description procedure is used to find correspondences from these standard images. Other works [65] rely on segmentations from ultrasound and magnetic resonance images to align them. Obtaining these segmentations may be costly, add additional computational complexity to the registration procedure, or be specific to the anatomies/modalities in question. In contrast, our method can be applied generally to any registration problem. In addition, we present a variant of our model which only relies on the images themselves during training. In our experiments, we find that this variant outperforms state-of-the-art baselines while also performing comparably to a variant of our model which leverages segmentations.
3 BrainMorph
Our tool, BrainMorph, is based on the previously published KeyMorph framework and we refer the reader to prior papers for more details [15, 16]. Let $(\bm{x}_{m},\bm{x}_{f})$ be a moving (source) and fixed (target) image [^1] Although we consider 3D volumes in this work, BrainMorph is agnostic to the number of dimensions. The terms “image” and “volume” are used interchangeably. pair, possibly of different contrasts or modalities. Additionally, we denote by $\mathcal{T}_{\theta}$ a parametric coordinate transformation with parameters $\theta$, such as those discussed in Appendix .1. The goal is to find the optimal transformation $\mathcal{T}_{\theta^{*}}$ such that the registered image $\bm{x}_{r}=\bm{x}_{m}\circ\mathcal{T}_{\theta^{*}}$ aligns with some fixed image $\bm{x}_{f}$, where $\circ$ denotes the spatial transformation of an image.
BrainMorph works by detecting $N$ keypoints $\bm{P}\in\mathbb{R}^{D\times N}$ from any given image. In this work, $D=3$. The keypoints are detected by a neural network $f_{w}$. Since $f_{w}$ detects keypoints for any image, the keypoints for any arbitrary image pair are guaranteed to be in correspondence by construction. Given corresponding keypoint sets $\bm{P}$ and $\bm{Q}$, the optimal transformation can be derived using a keypoint solver, which outputs the optimal transformation parameters as a function of the keypoints: $\theta^{*}(\bm{P},\bm{Q})$. Optionally, a vector of weights $\bm{w}$ can weight the correspondences, such that lower weights lead to a lower contribution to the overall alignment: $\theta^{*}(\bm{P},\bm{Q},\bm{w})$.
Fig. 1 depicts a graphical overview of BrainMorph. Note this formulation unlocks the benefits of keypoint-based registration, including robustness to large misalignments and interpretability (as compared to other learning-based methods) via visualizing the keypoints. Moreover, this formulation enables controllability in the sense that different transformations can be used to align the keypoints. In particular, during training, this allows for heterogeneity in training, such that the model can be robust to a wide variety of transformation types. At test-time, one can generate a dense set of registrations; the controllable nature of this framework enables the user to select the preferred registration.
3.1 Keypoint Detection Network
BrainMorph can leverage any deep learning-based keypoint detector [55, 59, 61]. In this work, we are interested in preserving translation equivariance; to this end, we leverage a center-of-mass (CoM) layer [66, 67] as the final layer, which computes the center-of-mass for each of the $N$ activation maps. This specialized layer is (approximately) translationally-equivariant and enables precise localization. Since the CoM layer expects positive values at every grid location, we insert a ReLU activation before the CoM layer.
3.2 Training
Training BrainMorph involves optimizing the learnable parameters within the CNN $f_{w}$ for pairwise registration. During training, we randomly sample pairs of moving and fixed images, and the general objective is:
$\begin{split}\operatorname*{arg\,max}_{w}\ &\mathbb{E}_{(\bm{x}_{m},\bm{x}_{f})}\ \mathcal{L}_{sim}\left(\bm{x}_{m}\circ\mathcal{T}_{\theta^{*}},\bm{x}_{f}\right)\\ &\text{where }\ \ \theta^{*}=\theta^{*}\left(f_{w}(\bm{x}_{f}),f_{w}(\bm{x}_{m}),\bm{w}\right)\end{split}$ (1)
where $\mathcal{L}_{sim}(\cdot,\cdot)$ measures image similarity between its two inputs. Weights $\bm{w}$ for keypoints correspondences are found as follows. First, we compute the energy (i.e. the aggregated sum) of each of the $N$ activation maps for both the fixed and moving image. Then, we multiply the corresponding energies, and compute the softmax of the $N$ energies to arrive at normalized weights [56].
In this work, we choose transformations whose optimal parameters can be solved in a closed-form and differentiable manner so that $f_{w}$ can be trained in an end-to-end fashion. Thus, the neural network is incentivized to detect $N$ anatomically-consistent keypoints from a given image, such that a good registration can be achieved. Note that we do not rely on any ground truth keypoints as supervision.
The BrainMorph framework enables flexibility in training depending on the choice of the loss function and the transformation used. $\mathcal{L}_{sim}$ can be any similarity function and can vary during training depending on the image pairs. In this work, we use MSE or Dice loss, depending on the current image pair. The closed-form optimal solution $\theta^{*}$ can depend on a hyperparameter $\lambda$, such as in TPS, which can be set to a constant or sampled from a distribution $\lambda\sim p(\lambda)$ during training. More details on the training details we used in this work are presented in Section 4.2. Once the model is trained, it can be used for both pairwise and groupwise registration during inference time, as described below.
3.3 Pairwise Registration
Pairwise follows straightforwardly from the training setup. Given a fixed image $\bm{x}_{f}$ and moving image $\bm{x}_{m}$, BrainMorph performs pairwise registration as follows:
$\begin{split}\bm{x}_{r}=\bm{x}_{m}\circ\mathcal{T}_{\theta^{*}}\ \text{where}\ \theta^{*}=\theta^{*}\left(f_{w}(\bm{x}_{f}),f_{w}(\bm{x}_{m})\right)\end{split}$ (2)
Note that at test time, we can use any transformation $\mathcal{T}_{\theta}$ (e.g, rigid, affine or TPS with any hyperparameter value), which would yield a different alignment based on the same keypoints.
3.4 Groupwise Registration
Groupwise registration methods try to mitigate uncertainties associated with any one image by simultaneously registering all images in a population. This incorporates all image information in the registration process and eliminates bias towards a chosen reference frame.
We propose a novel algorithm for groupwise registration based on detected keypoints. In the groupwise setting, we have $N$ subjects to align; thus we are solving for the optimal transformation as well as the optimal average space simultaneously. To achieve this with $\{\bm{P}_{i}\}_{i=1}^{N}$ keypoints, we optimize for the average space and the optimal transformations in an iterative, coordinate-ascent strategy by alternating the following two steps:
  1. 1.
    Given points $\bm{P}_{k}$ at the current iteration $k$, compute the average keypoints $\bar{\bm{P}}_{k+1}$
  2. 2.
    Compute new points $\bm{P}_{k+1}$ by registering all points $\bm{P}_{k}$ to $\bar{\bm{P}}_{k+1}$
After $K$ iterations, the algorithm converges to the average space $\bar{\bm{P}}_{K}$, and the final registration for the $i$’th image is found by transforming the image according to the transformation that aligns points $\bm{P}_{0}$ to $\bm{P}_{K}$.
Note that this algorithm only relies on keypoints. Thus, it is computationally efficient since keypoints can be precomputed and done serially, whereas other works must fit all images in memory at once. In our experiments, we demonstrate the scalability of our approach by registering more than 100 volumes simultaneously.
Figure 2: Moving, fixed, and aligned images in axial, sagittal, and coronal mid-slices. Keypoints extracted by BrainMorph are overlaid. For Aligned, aligned keypoints are dots and fixed keypoints are crosses. Color of the keypoint corresponds to depth. Note that keypoint weights are not visualized.
4 Materials and Methods
4.1 Dataset
We train BrainMorph on a massive dataset of over 100,000 images from nearly 16,000 unique subjects. All datasets are gathered from publicly-available brain studies, datasets, and challenges. The full list of datasets is given in the Appendix. Our tool requires the following two pre-processing steps for all image inputs: resampling to 1mm isotropic and cropping/padding to 256x256x256. Min-max rescaling to [0, 1] is performed as the first layer in the network, and thus we do not consider it a pre-processing step.
For purposes of training, we reorient all brains to MNI space  [^2] During training, we apply random affine transformations as an augmentation strategy. and perform skull-stripping with HD-BET [68], a robust deep learning-based skull-stripping tool. For images without extreme lesions, we further generate segmentations with SynthSeg [69, 70], which produces parcellations of 33 brain regions. We do not perform segmentation on images with extreme lesions. See the Appendix for the full list of brain regions.
4.2 Training Details
We are interested in learning foundational keypoints for the end goal of general-purpose brain registration. Thus, the keypoints should be optimized such that they are robust to a variety of brain MRI modalities and transformation types (rigid, affine, and nonlinear). Note that, BrainMorph is amenable to a variety of training strategies (pairwise sampling, loss function, and transformation type). We would like the network to be able to handle uni-modal, multi-modal, and longitudinal image pairs, with transform types including rigid, affine, and TPS. To do so, we perform heterogeneous training with different tasks for a single foundation model, where the task is randomly sampled in each mini-batch.
During training, we use two loss types: Dice of segmentation labels and mean-squared-error (MSE) of pixel values. We use three transformation types: rigid, affine, and Thin-plate splines, or TPS (nonlinear). TPS has a hyperparameter $\lambda$ which controls the degree of nonlinearity.
Table 2: Summary of BrainMorph training.
Image pair Transform type Loss
Normal TPS Dice
Skullstripped, lesion Affine MSE
Skullstripped, longitudinal Rigid MSE
In general, we are constrained by the following rules.
  1. 1.
    To use MSE loss, we must sample skull-stripped same-modality pairs.
  2. 2.
    For Dice loss, we may sample pairs which have corresponding segmentations (this precludes brains with lesions, for which SynthSeg [69, 70] cannot reliably segment).
  3. 3.
    For longitudinal image pairs, we use rigid transformation to simulate realistic downstream usage.
  4. 4.
    For image pairs with lesions, we use a restrictive affine transformation, as TPS will not guarantee bijective correspondence between images.
Table 2 summarizes the training strategy used, which we choose according to the above constraints. At every training iteration, we sample uniformly across the three image pair types. Thus, the model is trained to optimize the registration performance across all three image tasks with equal weighting.
We experimented with $N$ = 128, 256, and 512 keypoints, and perform a thorough analysis of the relationship between number of keypoints and registration performance in Section 6.2.2. For TPS transformations, we sample $\lambda$ during training from a log-uniform distribution $p(\lambda)=\text{LogUnif}(0,10)$. In addition, during training, in each mini-batch, we compute TPS on 32 keypoints chosen uniformly at random [71]. This is because TPS has a high memory requirement due to computing pairwise distances between every keypoint and grid location. Note that at test-time, we compute TPS on the full set of keypoints, and compute pairwise distances in a chunk-wise, iterative fashion to bypass this memory requirement. When minimizing Dice loss, we sample 14 regions uniformly at random for computational purposes.
For all models, we used a batch size of 1 image pair and the Adam optimizer [72] for training. We train for a total of 160K gradient steps. The following uniformly-sampled augmentations were applied to the moving image across all dimensions during training: rotations $[-180\degree,+180\degree]$, translations $[-30,30]$ voxels, scaling factor $[0.8,1.2]$, and shear $[-0.1,0.1]$. All training and GPU testing was performed on a machine equipped with an AMD EPYC 7513 32-Core processor and an Nvidia A100 GPU. CPU testing was performed on a machine equipped with an Intel Xeon Gold 6330 CPU @ 2.00GHz. All BrainMorph models are implemented in PyTorch.
4.3 Model Details
Table 3: BrainMorph backbone variants.
# parameters # downsampling layers
BrainMorph-S 4M 4
BrainMorph-M 16M 5
BrainMorph-L 66M 6
Our architecture backbone consists of a truncated UNet, which is identical to a standard UNet except all layers which operate at the original resolution (e.g. after the last upsampling layer) are removed [73]. All truncated UNets we use have two convolutional blocks at each resolution. Thus, the final center-of-mass layer extracts keypoints at half-resolution. This enables us to train deeper networks with a bottleneck operating on a very coarse grid, which we empirically find leads to better performance. In particular, we report results on three variants of the truncated UNet, which differ in the capacity as a function of the number of downsampling layers. We refer to them as BrainMorph-S, BrainMorph-M, and BrainMorph-L, and summarize them in Table 3. All references to BrainMorph are BrainMorph-L models, unless otherwise noted.
Model CPU Time, pre-processed GPU Time, pre-processed CPU Time, raw GPU Time, raw
ANTs, Rigid 101.38$\pm$2.33 - 143.38$\pm$2.33 -
ANTs, Affine 110.45$\pm$2.94 - 142.45$\pm$2.94 -
ANTs, Syn 216.03$\pm$3.14 - 248.03$\pm$3.14 -
BrainMorph, Rigid 109.84$\pm$1.80 1.05$\pm$0.29 108.27$\pm$1.82 1.05$\pm$0.25
BrainMorph, Affine 109.63$\pm$1.84 1.04$\pm$0.36 110.31$\pm$1.90 1.05$\pm$0.32
BrainMorph, TPS 180.14$\pm$1.99 1.24$\pm$0.30 180.40$\pm$1.91 1.25$\pm$0.31
Table 4: Average computation time in seconds for pairwise registration across different models. We separate between times when performing registration with pre-processed data and raw data. For ANTs, pre-processing requires skull-stripping with HD-BET, which adds an additional 32 seconds [68]. For SynthMorph, pre-processing requires an initial robust affine registration into a reference space. KeyMorph timings are based on KeyMorph-S.
4.4 Self-supervised Pretraining
We employ the following self-supervised pre-training strategy to aid in keypoint detector initialization, essentially encouraging equivariance of the keypoint extractor with respect to affine image deformations. Note that past works have leveraged equivariant networks [74]; however, we find these networks unstable to train and lack the capacity to capture variability present in our large datasets. Using a single subject, we pick a random set of keypoints $\bm{P}_{0}$ by sampling uniformly over the image coordinate grid. During pre-training, we apply random affine transformations to the input image as well as $\bm{P}_{0}$, and minimize the following keypoint loss:
$\operatorname*{arg\,min}_{w}\sum_{i}\mathbb{E}_{\bm{A}}\left\lVert\bm{A}\bm{P}_{0}-f_{w}\left(\bm{x}^{(i)}\circ\bm{A}\right)\right\rVert^{2}_{2}.$ (3)
Here, $\bm{A}$ is an affine transformation drawn from a uniform distribution over the parameter space.
We train for a total of 480K gradient steps. We use the same augmentation strategy as training, except that we linearly increase the degree of augmentation such that maximum augmentation is reached after 160K gradient steps. We use the same dataset for training and pretraining. Note that we assume that all the training data are in the same orientation and roughly in the center of the image in order for the sampled keypoints to apply well to all images in the dataset.
5 Experimental Setup
5.1 Evaluation Datasets
We use the following datasets for evaluation of all models. Note that these datasets are not included in training.
  1. 1.
    For healthy pairwise and cross-subject groupwise experiments, we evaluate on the IXI brain MRI dataset [^3] https://brain-development.org/ixi-dataset/. Each subject has T1, T2, and PD-weighted 3D MRI scans in spatial alignment, so we can use this dataset for both unimodal and multimodal registration experiments. We evaluate on 100 subjects.
  2. 2.
    For diseased subjects with lesions, we evaluate on the test split of the RSNA-ASNR-MICCAI BraTS dataset (to ensure no dataset contamination), which consists of adult brains with gliomas acquired with T1, T1gd, T2, and FLAIR sequences [^4] https://www.rsna.org/rsnai/ai-image-challenge/brain-tumor-ai-challenge-2021. We also evaluate on a dataset of brains with multiple sclerosis [75], which consists of 60 MS patients with T1, T2, and FLAIR sequences.
  3. 3.
    For longitudinal experiments, we evaluate on the OASIS2 dataset, which consists of longitudinal MRI brains in nondemented and demented older adults [^5] https://www.oasis-brains.org/. The number of timepoints per subject ranges between 2 and 10.
For all evaluation datasets, we perform resampling to 1mm isotropic and cropping/padding to $256^{3}$ image size.
5.2 Test-time Performance Evaluation
For pairwise experiments, we use each test subject as a moving volume $\bm{x}_{m}$, paired with another random test subject treated as a fixed volume $\bm{x}_{f}$. We simulate different degrees of misalignment by transforming $\bm{x}_{m}$ using rotation. Rotation is applied to all 3 axes at the specified degree. We use the predicted transformation to resample the moved segmentation labels on the fixed image grid. Unimodal/multimodal registration is an independent variable in our experiments. We experiment with rigid, affine, and nonlinear registration types for all models and baselines.
For longitudinal experiments, we perform groupwise registration on all available timepoints, and restrict to rigid transformations only. For cross-subject groupwise experiments, we sample different subject, same-modality images and experiment with varying group sizes in [4, 8, 16, 32, 64, 128]. We restrict to nonlinear transformations only. Similar to pairwise experiments, we simulate different degrees of misalignment by transforming all images using rotation applied to all 3 axes at the specified degree.
5.3 Metrics
For all experiments, we quantify alignment quality and properties of the transformation using Dice overlap score and Hausdorff distance (HD).
Figure 3: Dice performance on pairwise registration. Higher is better. Unimodal/multimodal, skull-stripped/non-skull-stripped.
Figure 4: HD performance on pairwise registration. Lower is better. Unimodal/multimodal, skull-stripped/non-skull-stripped.
5.4 Baselines
As the goal of this work is to develop a general-purpose tool for brain MRI registration, we wish to compare our proposed model against state-of-the-art and easily-accessible tools for this purpose. Our intended users are practitioners who desire a simple, easy-to-use tool that performs registration with minimal pre-processing or setup. Thus, we adhere to all instructions required by baselines (including intensity normalization, initial robust registration, etc.), and do not assume that data are pre-processed. Note that since skullstripping vs. non-skullstripping is an independent variable in our experiments, we do not perform any skullstripping as part of any baseline’s preprocessing requirements.
Figure 5: Pairwise registration results for BrainMorph and selected baselines. In first row, a 90 degree rotation is applied. In second row, a 135 degree rotation is applied.
  • ITK-Elastix is a widely-used software package which supports pairwise and groupwise registration [18]. Rigid, affine, and bspline alignments are supported. For all registrations, we perform a multi-resolution pyramid strategy at 4 resolutions in order to improve the capture range and robustness of the registration. Note that amongst all baselines, ITK-Elastix is the only one to provide a groupwise registration method. The method uses a 4D (3D+time) free-form B-spline deformation model and a similarity metric that minimizes variance of intensities under the constraint that the average deformation over images is zero. This constraint defines a true mean frame of reference that lie in the center of the population without having to calculate it explicitly. [^6] https://readthedocs.org/projects/simpleelastix/downloads/pdf/latest/
  • Advanced Normalizing Tools (ANTs) is a widely-used software package which is state-of-the-art for medical image registration [20]. We use the “Rigid” and “Affine” implementation for the rigid and affine model, respectively. The volumes are registered successively at three different resolutions: 0.25x, 0.5x and finally at full resolution. At 0.25x and 0.5x resolution, Gaussian smoothing with $\sigma$ of two and one voxels is applied, respectively. For non-linear registration, we use “SyN”, which performs Symmetric Normalization [17]. Finally, we used mutual information as the similarity metric for all models, which is suitable for registering images with different contrasts.
  • SynthMorph is a deep learning-based registration method which achieves agnosticism to modality/contrast by leveraging a generative strategy for synthesizing diverse images, thereby supporting multi-modal registration [14]. SynthMorph accepts as input the moving and fixed images and outputs a dense deformation field instead of global affine parameters, which is a common strategy in many well-performing registration models [13].
    Note that an important pre-processing step required by SynthMorph is an affine-registration step to a pre-defined reference space [76, 77]. Thus, this limits the flexibility of SynthMorph in the sense that all registrations are performed in this reference space. In contrast, BrainMorph enables the user to define any arbitrary reference space via the fixed image. In addition, this requirement increases pre-processing time. In our experiments, we first affine register every image using ANTs (see above). SynthMorph models are implemented in Keras/Tensorflow.
6 Results
6.1 Main Results
6.1.1 Pairwise Registration
We analyze the performance of baselines and our proposed BrainMorph under conditions of large initial misalignments in terms of rotation. Figs. 3 and 4 plot overall Dice and HD across rotation angle of the moving image for baselines and BrainMorph. Each panel depicts rigid, affine, and nonlinear registrations, respectively. Each separate figure varies different combinations of unimodal vs. multimodal, skull-stripped and non-skull-stripped.
We find that all baseline models suffer substantially as the rotation angle increases, across all transformation types. BrainMorph performs well across all rotation angles, and is particularly strong for rigid and affine transformations. ANTs and SynthMorph yield excellent results when the initial misalignment is small (e.g. near 0 degrees of rotation) and when data is skull-stripped, particularly for nonlinear transformations. However, the accuracy drops substantially when the misalignment is not minimal.
We compare the computational time across different models in Table 4. Some representative examples for all models is provided in Fig. 5 for qualitative evaluation. Overall, BrainMorph outperforms other baselines at high degrees of initial misalignment, and furthermore performs comparably or often better (at large misalignments) than the state-of-the-art ANTs registration, while requiring substantially less runtime.
6.1.2 Groupwise Registration
Figure 6: Groupwise registration with group size of 4. For ITK-Elastix, B-spline registration is used. For BrainMorph, TPS with $\lambda=0$ is used. The last column shows the average brain in the optimized template space.
Figure 7: Groupwise registration using BrainMorph and ITK-Elastix. For BrainMorph, the transformation used is TPS with $\lambda=0$. For ITK-Elastix, the transformation is nonlinear b-spline. We observe significant improvement in groupwise registration capabilities of BrainMorph, at a time reduction of nearly 4 orders of magnitude.
Fig. 7 depicts groupwise performance and timings for BrainMorph vs. ITK-Elastix. We use B-spline for ITK-Elastix and TPS with $\lambda=0$ for BrainMorph. We find that BrainMorph has much better and more stable performance across all group sizes we tested. Note that on our CPU, ITK-Elastix failed on 128 subjects. BrainMorph has an advantage in that keypoints can be precomputed in a serial fashion on a GPU, thus enabling much better scaling to large group sizes.
Fig. 6 depicts a representative example of groupwise registration with 4 subjects. The first row depicts the initial unaligned images, and the second and third row show the results of ITK-Elastix and BrainMorph, respectively. The last column depicts the average template brain for all 4 subjects. We find that BrainMorph groupwise registration to be substantially better, as evidenced by the sharp lines in the template brain. On a GPU, BrainMorph is also faster than ITK-Elastix by nearly 4 orders of magnitude.
6.1.3 Longitudinal Registration
Figure 8: a) Registration performance on longitudinal registration. Rigid transformations for all models. b) Registration performance on subjects with lesions. Affine transformations for all models, except for SynthMorph which only supports dense.
Fig 8(a) shows boxplots for longitudinal registration performance across rotation angles for skull-stripped and non-skull-stripped, respectively. We find that across all rotation angles and with and without skullstripping, BrainMorph outperforms ITK-Elastix.
6.1.4 Registration with Lesions
Fig 8(b) shows a boxplot of performance of all models on lesion data. We observe that the weighted variant of BrainMorph outperforms all baselines across most rotation angles and is generally has more stable performance. In particular, weighted BrainMorph tends to outperform the unweighted variant of BrainMorph at 45 degrees of rotation and above.
Note that we use Dice performance as a proxy for registration quality, even though the SynthSeg-generated segmentations are not guaranteed to be robust to diseased patients.
6.2 Keypoint Analysis
6.2.1 Visualizing keypoints
In contrast to existing models that compute the transformation parameters using a “black-box” neural network, one can investigate the keypoints that BrainMorph learns to drive the alignment. Fig. 2 shows the keypoints for a moving and fixed subject pair via mid-slices for sagittal, axial, and coronal views. The first three columns depict keypoints extracted from skull-stripped images, and the last three columns depict keypoints extracted from non-skull-stripped images. The color of the keypoints represents depth with respect to the mid-slice. The “Aligned” slices show both warped (dots) and fixed (crosses) points.
Note that keypoint locations are trained end-to-end without explicit annotations. We observe that keypoint locations are generally in sub-cortical regions, where anatomical variability is relatively low across subjects as compared to cortical regions.
6.2.2 Number of Keypoints
Figure 9: Registration performance of variants of BrainMorph and varying the number of keypoints.
As an ablation, we examine the effect of the number of keypoints used for alignment across different transformations. We trained BrainMorph model variants with 128, 256, and 512 keypoints. Fig. 9 illustrates that performance is not discernably correlated to increasing the number of keypoints. We hypothesize that as keypoints are generally in subcortical regions which are anatomically stable, increasing the number of correspondences does not provide further advantage beyond a certain point.
7 Discussion
The results demonstrate that BrainMorph is a robust and flexible tool for brain MRI registration. On pairwise registration, BrainMorph is generally superior to baselines across all degrees of initial misalignment, and is state-of-the-art for affine and rigid transformations. These results hold consistently for for unimodal and multimodal registration, as well as skull-stripped and non-skull-stripped data. In addition, BrainMorph does not require extensive pre-processing like skullstripping and pre-affine registration. On longitudinal and groupwise registration, BrainMorph is superior to baseline registration algorithms, while being much more memory efficient and nearly 4 orders of magnitude faster.
The main limitation of BrainMorph is nonlinear performance at low initial misalignment and skull-stripped data, for which ANTs and SynthMorph perform excellently. For this reason, users who require good nonlinear registrations may consider using BrainMorph as a robust initial alignment tool, and further performing nonlinear registration using a tool like ANTs.
8 Conclusion
We presented a robust and flexible registration tool based on the KeyMorph framework, called BrainMorph, which is a deep learning-based image registration method that uses corresponding keypoints to derive the optimal transformation that align the images. This formulation enables interpretability, robustness to large initial misalignments, and flexibility/controllability of registrations at test-time. Training on a massive dataset of over 100,000 unique images from nearly 16,000 subjects enables our tool to work on raw data with minimal pre-processing. Empirically, we demonstrate fast, state-of-the-art performance across rigid, affine, nonlinear, and groupwise registration, particularly at large degrees of initial misalignment.